State the first term, and write the next four terms of the sequences;
a) 9; 14; 19
b) 3; 7; 11
c) 8;15; 22
d) 4; 11; 18
1
Expert's answer
2020-07-13T18:50:11-0400

a)In this arithmetic sequence, the first term is 9. The common difference d

 is 5. You can find this by subtracting a term from the consecutive term: 14 - 9 = 5

In order to find the next four terms, keep adding 5 to the previous term:

19+5=24

24+5=29

29+5=34

34+5=39

b)In this arithmetic sequence, the first term is 3. The common difference d

 is 4. You can find this by subtracting a term from the consecutive term: 7 - 3 = 4

In order to find the next four terms, keep adding 4 to the previous term:

11+4=15

15+4=19

19+4=23

23+4=27

c)In this arithmetic sequence, the first term is 8. The common difference d

 is 7. You can find this by subtracting a term from the consecutive term: 15 - 8 = 7

In order to find the next four terms, keep adding 7 to the previous term:

22+7=29

29+7=36

36+7=43

43+7=50

d)In this arithmetic sequence, the first term is 4. The common difference d

 is 7. You can find this by subtracting a term from the consecutive term: 11 - 4 = 7

In order to find the next four terms, keep adding 7 to the previous term:

18+7=25

25+7=32

32+7=39

39+7=46